### v1

# Diophantine Graphs

Abstract
This manuscript introduces Diophantine labeling, a new way of labeling of the vertices for finite simple undirected graphs with some divisibility condition on the edges. Maximal graphs admitting Diophantine labeling are investigated and their number of edges are computed. Some number-theoretic techniques are used to characterize vertices of maximum degree and nonadjacent vertices. Some necessary and sufficient conditions for vertices of equal degrees are found.
Keywords: Graph labeling, Prime graph, Diophantine graph, $p$-adic valuation.
MSC code: 05A10, 05C07, 05C78, 11A05, 11A25, 11B75, 11D04, 11D88.
1 Introduction
A graph labeling is an assignment of real values to the vertices or edges that meet some conditions. The definition of prime labeling was originated by R. Entringer and introduced formally in a paper by R. Tout [12]. A graph $G=(V,E)$ with $n$ vertices has a prime labeling if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))=1$. Many results of prime labeling and other types of prime labeling can be found in the dynamic survey of graph labeling by J.A. Gallian [6]. Other variants of prime labeling have been introduced by many researchers (see for example [4], [5], [8], [13]).
This paper gives a generalization for the concept of prime graphs. More precisely, we define a graph $G$ with $n$ vertices to be a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))\mid n$. Clearly, any prime graph is a Diophantine graph. Moreover, any Diophantine graph with a prime number of vertices is a prime graph. In the maximal Diophantine graph with $n$ vertices, an explicit formulas of the number of edges and the degree of a vertex are proved, the full degree vertices are determined. Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among non-full degree vertices in the maximal Diophantine graphs.
The paper is organized as follows. Section 2 includes some definitions, preliminaries, lemmas and direct observations. Section 3 presents the number of edges formula in the maximal Diophantine graph with $n$ vertices. Section 4 is partitioned into three subsections, each unveiling important results related to the maximal Diophantine graph with order $n$. Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that compute the degrees of any vertex. Subsection 4.2 and 4.3 encompass some necessary and sufficient conditions for equality of degrees.
For definitions and terminologies of graph theory, we follow F. Harary [7] and A. Bickle [2]. Also, we follow D. Burton [3] and K.H. Rosen [9] for basic definitions and notations in number theory.
2 Definitions and Preliminaries
In the beginning of this section, we give the definitions of the prime graphs and Diophantine graphs. It also explores the relation between them and some direct observations.
Definition 2.1.
[11], [12] Let $G$ be a graph with $n$ vertices. The graph $G$ is called a prime graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))=1$. Such a map $f$ is called a prime labeling of $G$. The maximal prime graph with $n$ vertices, denoted by $R_{n}$, is a prime graph such that adding any new edge yields a non-prime graph.
Definition 2.2.
Let $G$ be a graph with $n$ vertices. The graph $G$ is called a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))\mid n$. Such a map $f$ is called a Diophantine labeling of $G$. The maximal Diophantine graph with $n$ vertices, denoted by $D_{n}$, is a Diophantine graph such that adding any new edge yields a non-Diophantine graph.
Example 2.1.
The Petersen graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
66662222333344441010101011117777888899995555
11112222333344445555
Figure 1: $P$ is a Diophantine graph while $K_{5}$ is not
Example 2.2.
The following are some maximal Diophantine graphs.
1111222233334444
11112222333344445555
111122223333444455556666
1111555577773333222266664444
11112222333344445555666677778888
Figure 2: Maximal Diophantine Graphs $D_{4},\dots,D_{8}$
It is easy to see that every prime graph is a Diophantine graph, but the converse is not necessarily true, e.g., $D_{4},D_{6},D_{8}$. However, in the case that $n$ is a prime number, a graph is Diophantine if and only if it is a prime graph. Moreover, $D_{n}\cong R_{n}.$ The proof is straightforward using the following lemma.
Lemma 2.1.
Let $n>1$. Then we have $n$ is a prime number  if and only if  for all $a,b\leq n$, $a\neq b$, [$(a,b)=1$ if and only if $(a,b)\mid n$].
Proof.
Let $n$ be a prime number, $a,b<n$ and $a\neq b$. If $(a,b)=1,$ then $(a,b)\mid n$. On the other hand, if $(a,b)\mid n,$ then either $(a,b)=1$ or $(a,b)=n$. In the case that $(a,b)=1$, we have nothing to proof. In the other case that $(a,b)=n,$ we get $a=b=n$ which is a contradiction with the hypothesis.
Conversely, Assume $a\neq b$ and [$(a,b)=1$ if and only if $(a,b)\mid n$] for all $a,b<n$. Suppose, by the way of contradiction, $n>1$ is a composite number. Then there exists prime $p\mid n$ such that $p\leq\frac{n}{2}$. Let $a=p\leq n$, $b=2p\leq n$. Hence, $(a,b)=p\mid n$ and $(a,b)\neq 1,$ which contradicts the assumption. Hence, the result follows. ∎
In the subsequent sections, in any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and we denote the set of all multiples of a positive integer $a$ up to $n$ by $M_{a}$, i.e. $M_{a}:=\{ta\ :\ t=1,2,\dots,\lfloor\frac{n}{a}\rfloor\}.$ Now, we present a crucial definition and an important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
Definition 2.3.
For a given positive integer $n$ and prime number $p$, the successor of the $p$-adic valuation is defined by $\acute{v}_{p}(n):=v_{p}(n)+1,$ where $v_{p}(n)$ is the $p$-adic valuation of an integer $n.$ The number $p^{\acute{v}_{p}(n)}$ is called the critical prime power number with respect to $p,n$.
Lemma 2.2.
Let $u,v\in V(D_{n})$ and $f$ be a Diophantine labeling of $D_{n}$. Then we have
  • (i)
    for every prime number $p$, $p\mid f^{*}(u)$ if and only if $p^{\acute{v}_{p}(n)}\mid f(u);$
  • (ii)
    if $f(u)\mid f(v)$, then $f^{*}(u)\mid f^{*}(v).$
Proof.
The proof of part $(i)$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
$v_{p}\left(\frac{b}{a}\right):=v_{p}(b)-v_{p}(a),$
for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p.$ Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
$\begin{split}v_{p}\left(\frac{f(v)}{f(u)}\frac{(f(u),n)}{(f(v),n)}\right)&=v_{p}(f(v))-v_{p}(f(u))+v_{p}\big{(}(f(u),n)\big{)}-v_{p}\big{(}(f(v),n)\big{)}.\\ &=v_{p}(f(v))-v_{p}(f(u))+\min\big{\{}v_{p}(f(u)),v_{p}(n)\big{\}}-\min\big{\{}v_{p}(f(v)),v_{p}(n)\big{\}}.\end{split}$
Therefore, one can easily check that
$v_{p}\left(\frac{f(v)}{f(u)}\frac{(f(u),n)}{(f(v),n)}\right)\geq 0$
in all possible cases between $v_{p}(f(u)),v_{p}(n)$ and $v_{p}(f(v))$, namely
$v_{p}(n)<v_{p}(f(u)),\quad v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\quad\mbox{and}\quad v_{p}(f(v))<v_{p}(n).$
Hence,
$\frac{f(u)}{(f(u),n)}\ \Bigg{|}\ \frac{f(v)}{(f(v),n)},$
which completes the proof. ∎
In the proof of Lemma 2.3, we will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$.
Lemma 2.3.
For every $n\in\mathbb{Z}^{+}$, there exists $j\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\ldots,n\right\}$ such that
$\frac{n}{j+1}<\lfloor\sqrt{n}\rfloor\leq\frac{n}{j}\quad\mbox{and}\quad\frac{n}{j}-\frac{n}{j+1}<1.$
Proof.
Clearly, we have
$\begin{cases}\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}<1,&\quad\mbox{if}\quad n<\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1);\\ &\\ \lfloor\sqrt{n}\rfloor\notin\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}\geq 1,&\quad\mbox{if}\quad n\geq\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1).\\ \end{cases}$
Since $\lfloor\sqrt{n}\rfloor<\frac{n}{\lfloor\sqrt{n}\rfloor}$ and $\frac{n}{j}-\frac{n}{j+1}>\frac{n}{j+1}-\frac{n}{j+2},$ therefore there exists $j\geq\lfloor\sqrt{n}\rfloor$ such that
$\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{j+2},\frac{n}{j+1}\right]\quad\mbox{and}\quad\frac{n}{j+1}-\frac{n}{j+2}\leq\frac{n}{\lfloor\sqrt{n}\rfloor+1}-\frac{n}{\lfloor\sqrt{n}\rfloor+2}<1,$
which complete the proof. ∎
Lemma 2.4.
Let $a,b$ and $n\in\mathbb{Z}^{+}$. If $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$ and $a\neq b$, then $ab>n$.
Proof.
Let $a,b$ and $n\in\mathbb{N}$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$. Therefore, there exists $j\in\{1,2,\dots,n\}$ such that $\frac{n}{j+1}<a,b\leq\frac{n}{j}$. Consequently, $\frac{n}{j}-\frac{n}{j+1}>1$, since $a\neq b$. Using Lemma 2.3, we get that there exists $i\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\dots,n\right\}$ such that
$\frac{n}{i+1}<\lfloor\sqrt{n}\rfloor\leq\frac{n}{i}\quad\mbox{and}\quad\frac{n}{i}-\frac{n}{i+1}<1.$
Thus, $a,b>\lfloor\sqrt{n}\rfloor$. Hence, $ab>n$. ∎
Lemma 2.5.
Let $a,b,n\in\mathbb{Z}^{+}$ and let $p_{1},p_{2},q_{1},q_{2}$ be distinct prime numbers and $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ be positive integers. If $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}}$ such that
$\frac{n}{i+1}<p_{1}^{\acute{v}_{p_{1}}(n)},q_{1}^{\acute{v}_{q_{1}}(n)}\leq\frac{n}{i}\quad\mbox{and}\quad\frac{n}{j+1}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}\leq\frac{n}{j}$ (1)
for some $i,j=1,2,\dots,n$, then
$a>n\quad\mbox{or}\quad b>n.$
Proof.
Let $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}}$ such that $p_{1},p_{2},q_{1},q_{2}$ are distinct prime numbers satisfying equation (1), where $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}>0$ are positive integers. Consequently,
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$ (2)
Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.2 part $(i)$ implies that
$p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\ \Big{|}\ a\quad\mbox{and}\quad q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\ \Big{|}\ b.$ (3)
In the case $i=j,$ equation (1) implies that
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$
Therefore, using Lemma 2.4 and equation (3), we get
$n<p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\leq a\quad\mbox{and}\quad n<q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\leq b.$
If $i>j$, then equation (3) implies that
$p_{1}^{\acute{v}_{p_{1}}(n)}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}\quad\mbox{and}\quad q_{1}^{\acute{v}_{q_{1}}(n)}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}.$
Therefore, using Lemma 2.4 and equations (2) and (3), we get
$n<p_{1}^{\acute{v}_{p_{1}}(n)}q_{1}^{\acute{v}_{q_{1}}(n)}<p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\leq a.$
In a very similar way, one can prove that, if $i<j$, then
$n<p_{2}^{\acute{v}_{p_{2}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}<q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\leq b.$
Hence, the result follows. ∎
Lemma 2.6.
Let $a,b,t,n\in\mathbb{Z}^{+}$. If $\left\lfloor\frac{n}{a}\right\rfloor=\left\lfloor\frac{n}{b}\right\rfloor$, then $\left\lfloor\frac{n}{ta}\right\rfloor=\left\lfloor\frac{n}{tb}\right\rfloor$.
The proof of Lemma 2.6 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+}.$ Lemmas 2.5 and 2.6 will be used in the proof of Theorem 4.5.
3 Number of Edges of $D_{n}$
The following lemma is a characterization of nonadjacent vertices in the maximal Diophantine graphs $D_{n}$.
Lemma 3.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u,v\in V(D_{n}),$ $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that
$f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}}:=\left\{kp^{\acute{v}_{p}(n)}:\ k=1,2,\dots,\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\right\}.$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u,v\in V(D_{n})$. Then we have
$\begin{split}uv\notin E(D_{n})&\quad\mbox{if and only if}\quad(f(u),f(v))\nmid n\\ &\quad\mbox{if and only if}\quad v_{p}\left(\big{(}f(u),f(v)\big{)}\right)>v_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad\min\left\{v_{p}(f(u)),\ v_{p}(f(v))\right\}\geq v_{p}(n)+1=\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad v_{p}(f(u))\geq\acute{v}_{p}(n)\quad\mbox{and}\quad v_{p}(f(v))\geq\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad p^{\acute{v}_{p}(n)}\mid f(v)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad f(u),\ f(v)\in M_{p^{\acute{v}_{p}(n)}}\quad\mbox{for some prime number $p$.}\hskip 139.41832pt\square\end{split}$
It is clear that from the Lemma 3.1, if such a prime number $p$ exists, then $p<\frac{n}{2}$. Any edge in $D_{n}$ represents a solvable linear Diophantine equation $ax+by=n$ and any non-edge in $D_{n}$ represents an unsolvable linear Diophantine equation $ax+by=n$, more precisely, using lemma 3.1, we have for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if the linear equation $f(u)x+f(v)y=n$ is unsolvable.
Theorem 3.1.
The number of edges of the maximal Diophantine graph $D_{n}$ with order $n$ is given by
$C(n,2)-\sum_{1\leq i\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},2\Bigg{)}+\sum_{1\leq i<j\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg{\rfloor},2\Bigg{)}-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),$
where $p_{1},p_{2},\dots,p_{\pi(n)}<n$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient with $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
Proof. Let $D_{n}$ be the maximal Diophantine graph with $n$ vertices, $f$ be a Diophantine labeling of $D_{n}$ and $p_{i}<n,\ i=1,2,\dots,\pi(n)$ be distinct prime numbers. Using Lemma 3.1, we get for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$. Define
$S_{p_{i}}:=\big{\{}\{u,v\}\subset V(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}},\ f(u)\neq f(v)\big{\}},$
$i=1,2,\dots,\pi(n)$. So, we have the following equation
$\begin{split}|E(D_{n})|&=|E(K_{n})|-|E(\overline{D_{n}})|,\ \mbox{where $\overline{D_{n}}$ is the complement of $D_{n}$.}\\ &=|E(K_{n})|-\left|\bigcup\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|.\end{split}$
Using the inclusion-exclusion principle (see [10]), we have
$|E(D_{n})|=C(n,2)-\left[\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|-\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|+\dots+(-1)^{\pi(n)-1}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|\right].$
Therefore,
$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\end{split}$
4 Degrees of Vertices
4.1 The Degree of a Vertex in $D_{n}$
Theorem 4.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u\in V(D_{n})$, $p$ is a prime number,
$\deg(u)=n-1\quad\mbox{if and only if}\quad f(u)\mid n\quad\mbox{{or}}\quad\frac{n}{2}<f(u)=p^{\acute{v}_{p}(n)}<n,$
where the exclusive or will be typed in bold while the inclusive or is typed as usual.
Proof.
Assume that $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
$f(u)\nmid n\quad\mbox{and}\quad\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad\frac{n}{2}\nless f(u)=p^{\acute{v}_{p}(n)}\nless n\right),$
where $p$ is a prime number. In case of $f(u)=p^{\acute{v}_{p}(n)}$ and $1<f(u)<\frac{n}{2}$, we have
$(f(u),f(v))=(f(u),2f(u))=\left(p^{\acute{v}_{p}(n)},2p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
where $v\in V(D_{n})$ such that $f(v)=2f(n)<n$, which contradicts the full degree of $u$. In the other case, when $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we have that there exists a prime number $p\mid f(u)$ such that $v_{p}(f(u))>v_{p}(n).$ Then we get
$v_{p}\left(f(u)\right)\geq v_{p}(n)+1=\acute{v}_{p}(n),$
so $p^{\acute{v}_{p}(n)}\mid f(u)$. Therefore, one can see that
$\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
which is a contradiction of the full degree of $u$. Therefore, one can see that
$f(u)\nmid n\quad\mbox{{or}}\quad f(u)=p^{\acute{v}_{p}(n)},\quad\frac{n}{2}<p^{\acute{v}_{p}(n)}<n.$
Conversely, in the first case of $f(u)\mid n$, we have for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get any proper multiple of $f(u)$ is bigger than $n$. Since $f(u)=p^{\acute{v}_{p}(n)}$, therefore for all $v\in V(D_{n})$, $(f(u),f(v))\mid p^{v_{p}(n)}$, which means for all $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the both cases, we get $\deg(u)=n-1$. ∎
Similar results can be deduced in the case of the maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1. The properties of the maximal Diophantine graphs $D_{n}$ given in Lemma 3.1 and Theorem 4.1 are essential in what follows.
Corollary 4.1.
There are infinitely many Diophantine graphs which are not prime graphs.
Proof.
Let $D_{p^{k}}$ be the maximal Diophantine graphs with $p^{k}$ vertices, where $k>1$ and $p$ is a prime number. $D_{p^{k}}$ can not be a prime graph because the number full degree vertices in $R_{p^{k}}$ that is equal to
[…]
is less than of the number full degree vertices in $D_{p^{k}}$ which is at least
[…]
Hence, the proof follows. ∎
Lemma 4.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $f^{*}(u)=p^{k}$, for some $p$ is a prime number and $k\in\mathbb{N}$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$k=0.$}\\ n-\Big{\lfloor}\frac{n}{p^{\acute{v}_{p}(n)}}\Big{\rfloor},&\hbox{$k>0.$}\end{array}\right.$
Proof.
Let $D_{n}$ be the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u\in V(D_{n})$ such that $f^{*}(u)=p^{k}$, where $p$ is a prime number and $k\geq 0$. In the case of $k=0$, we have $f^{*}(u)=1$. Since $f^{*}(u)=\frac{f(u)}{(f(u),n)}$, we have $f(u)\mid n$. Therefore, using Theorem 4.1, $\deg(u)=n-1$. In the other case of $k>0$, we have
[…]
Let $s\in V(D_{n})-\{u\}$ such that
[…]
Therefore,
$\deg(u)=n-1-\Big{(}\Big{|}M_{p^{\acute{v}_{p}(n)}}\Big{|}-1\Big{)}=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
Hence, we have the proof. ∎
Example 4.1.
In $D_{20}$, let $f(u)=18$ be the label of a vertex $u\in D_{20}$. Then $f^{*}(u)=9=3^{2}$ and we have $k=2>0$. Therefore, Using Lemma 4.1, we get the following
$\deg(u)=20-\left\lfloor\frac{20}{3^{\acute{v}_{3}(20)}}\right\rfloor=14.$
Another way to compute the degree of $u$ with label $18$. The neighborhood of $u$ is the set of adjacent vertices of $u$ that has labels belong to the set $\{1,2,4,5,7,8,10,11,13,14,16,17,19,20\}$. Therefore, $\deg(u)=14$.
In a very similar way of Lemma 4.1, using the inclusion-exclusion principle and the identity $\sum\limits_{i=0}^{r}(-1)^{r}\binom{r}{i}=0$, one can prove the following theorem.
Theorem 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph with order $n$, $f$ is a Diophantine labeling and $u\in V(D_{n})$. If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\geq 0$, $i=1,2,\dots,r$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\bigg{\rfloor}+\sum\limits_{1\leq i.j\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\bigg{\rfloor}-\dots+(-1)^{r}\Bigg{\lfloor}\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},&\hbox{$f^{*}(u)>1.$}\end{array}\right.$
Similar results of Theorems 3.1, 4.2 can be obtained in the maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
4.2 Vertices of Equal Degrees $<n-1$
Lemma 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$. If $f(u)\mid f(v)$, then $N(u)\supseteq N(v)$, where $N(s)$ defines the neighborhood of $s$ as the set of all vertices in $D_{n}$ that join the vertex $s$.
Proof.
The proof is straightforward. ∎
Lemma 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)=N(v)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that
$f(u)\mid f(v)\quad\mbox{and}\quad\deg(u)=\deg(v).$
Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$. Then, using Lemma 4.2, $su\in E(D_{n})$. Hence, $N(u)\supseteq N(v)$. On the other hand, let $s\in V(D_{n})-\{u,v\}$ such that $su\in E(D_{n})$. Suppose by contrary that $sv\notin E(D_{n})$. Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)\supset N(v)$. Then we have $\deg(u)>\deg(v)$, which contradicts the hypothesis of $\deg(u)=\deg(v)$. Therefore, $sv\in E(D_{n})$. Consequently, $N(u)\subseteq N(v)$. Hence, $N(u)=N(v)$. ∎
It is clear that, if $N(u)=N(v)$, then $\deg(u)=\deg(v)$. Thus, under the assumption that $f(u)\mid f(v)$, we have
$N(u)=N(v)\quad\mbox{ if and only if}\quad\deg(u)=\deg(v).$
Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
Example 4.2.
In $D_{20}$, let $f(u)=7$ and $f(v)=8$, where $u,v\in V(D_{20})$. $N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$. Then we have $\deg(u)=\deg(v)$ while $f(u)\nmid f(v)$, but $N(u)\neq N(v)$.
Lemma 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $deg(u)<n-1$, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u\in V(D_{n})$ with $deg(u)<n-1$. Then there exists a vertex $s\in V(D_{n})$ such that $su\notin E(D_{n})$. Therefore, using Lemma 3.1, there exists a prime number $p$ such that $f(s),f(u)\in M_{p^{\acute{v}_{p}(n)}}$. Hence, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$. ∎
Theorem 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\geq 0$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are two prime numbers. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq t,\ \acute{v}_{q}(n)\leq k\quad\mbox{and}\quad\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k>0$, where $p,q$ are primes numbers and $u$ $v\in V(D_{n})$ with $deg(u)=deg(v)$, $deg(u)<n-1$, $deg(v)<n-1$. Therefore, using Lemma 4.4, we have
$p^{\acute{v}_{p}(n)}\mid f(u)=p^{t}\quad\mbox{and}\quad q^{\acute{v}_{q}(n)}\mid f(v)=q^{k}.$
Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. By using Lemma 4.1, we get
$deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Hence, we obtain the following equation
$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$. Then we get $p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{and}\ q^{\acute{v}_{q}(n)}\mid f(v).$ Thus, using Lemma 4.1, we have the following two equations
$\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Hence, the result follows. ∎
Corollary 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices with $\deg(u)<n-1$, $\deg(v)<n-1$ such that $f(u)=p^{t}$, $f(v)=p^{k}$ for some $t,k\geq 0$, where $p$ is a prime number. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq\min\{t,k\}.$
Theorem 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices such that $f(v)=tf(u)$ for some $t\geq 1$ and $f(v)$ is not a prime power. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if for every prime number}\ p\mid t,\quad\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}((f(v))).$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $f(v)$ is not a prime power. Proof by contrapositive, suppose that there exists a prime number $p_{0}\mid t$ such that
$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}((f(v))),$
or
$\acute{v}_{p_{0}}(n)>v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$
The first case contradicts the assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
$v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$ (4)
Let a vertex $s\in V(D_{n})$ such that $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Therfore, using equation (4), $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}$. Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$. Consequently, $s\neq u$. Also, $s\neq v$ for $f(v)$ is not a prime power. Moreover, using equation (4), we get
$(f(u),f(s))\mid n\quad\mbox{while}\quad(f(v),f(s))\nmid n.$
Thus,
$su\in E(D_{n})\quad\mbox{while}\quad sv\notin E(D_{n}).$
Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)\supset N(v)$. Hence, $\deg(u)\neq\deg(v)$.
Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$. Then, using Lemma 4.2, $N(u)\supset N(v)$. So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that
$s\in N(u)\quad\mbox{and}\quad s\notin N(v).$
Consequently,
$(f(u),f(s))\mid n\quad\mbox{and}\quad(f(v),f(s))\nmid n.$ (5)
Then, using Lemma 3.1, there exists a prime number $p_{0}$ such that
$f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}\quad\mbox{and}\quad p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u).$ (6)
Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation 6, we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation 5, we have the following
$v_{p_{0}}\left((f(u),f(s))\right)\leq v_{p_{0}}(n)\quad\mbox{and}\quad v_{p_{0}}\left((f(v),f(s))\right)>v_{p_{0}}(n).$ (7)
Hence, using equation 7, we get $v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)$ and $v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$ for some prime number $p_{0}\mid t$. ∎
4.3 More Necessary and Sufficient Conditions for Equality of Degrees
The following theorem is a generalization of one direction of Theorems 4.3.
Theorem 4.5.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that $\deg(u)<n-1$, $\deg(v)<n-1$. If $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}$ and for every prime number $p\mid f^{*}(u)$, there exits a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of prime factors of $n$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that
$\deg(u)<n-1,\quad\deg(v)<n-1\quad\mbox{and}\quad\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}.$
Therefore, $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}>0.$ Hence, there is a prime number $p$ such that $p\mid f^{*}(u)$. Consequently, our hypothesis implies that there is a prime number $q\mid f^{*}(v)$ such that
$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$ (8)
Therefore, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v).$ Then, using the fundamental theorem of arithmetic, there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \Big{|}\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \Big{|}\ f(v).$
By definition of $f(\grave{u})$ and $f(\grave{v})$, we have $f^{*}(u),f^{*}(\grave{u})$ have the same prime factors and $f^{*}(v),f^{*}(\grave{v})$ have the same prime factors. Since $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)},$ therefore
$\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$ (9)
Moreover, for every prime number $p$, we have
$p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{if and only if}\ p^{\acute{v}_{p}(n)}\mid f(\grave{u}).$
Therefore,
$\bigcup\limits_{p|f^{*}(u)}M_{p^{\acute{v}_{p}(n)}}=\bigcup\limits_{p|f^{*}(\grave{u})}M_{p^{\acute{v}_{p}(n)}}.$
Consequently, using Lemma 3.1, for each vertex $s\in V(D_{n})-\{u,\ \grave{u}\}$, we have
$su\in E(D_{n})\quad\mbox{if and only if}\quad s\grave{u}\in E(D_{n}).$
Hence,
$\deg(u)=\deg(\grave{u}).$ (10)
Similarly, we get
$\deg(v)=\deg(\grave{v}).$ (11)
Let $d=(f(\grave{u}),f(\grave{v}))$. Thus, Lemmas 2.2 part $(ii)$ implies that
$\frac{d}{(d,n)}\ \bigg{|}\ f^{*}(\grave{u}),f^{*}(\grave{v}),$ (12)
and hence
$\omega\left(\frac{d}{(d,n)}\right)\leq\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$
Let us consider the following cases.
Case 1: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right).$ Then, using equation (12), $\frac{d}{(d,n)}$, $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ have the same prime factors. A similar argument used to prove equation (10) implies that
$\deg(\grave{u})=\deg(\grave{v}).$
Hence, using equations (10) and (11), we get
$\deg(u)=\deg(v).$
Case 2: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+1=r.$ Then we get $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly one prime factor. Therefore, using the fundamental theorem of arithmetic, we obtain the following two equations
$f(\grave{u})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)},$
where $h_{i},p$ and $q$ are distinct prime numbers such that $h_{i}\mid d,p\nmid d$ and $q\nmid d$ for all $i=1,2,\dots,r-1$. Then, using Theorem 4.2, we have the following two equations
$\begin{split}\deg(\grave{u})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (13)
and
$\begin{split}\deg(\grave{v})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (14)
Using Lemma 2.6 and equations (8), (13) and (14), we get
$\deg(\grave{u})=\deg(\grave{v}),$
and hence
$\deg(u)=\deg(v).$
Case 3: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+2=r.$ Then $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly two prime factors. Therefore, using the fundamental theorem of arithmetic, we have the following two equations
$f(\grave{u})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)},$
where $h_{i},p_{1},p_{2},q_{1}$ and $q_{2}$ are distinct prime numbers satisfying $h_{i}\mid d$, $p_{1}\nmid d$, $p_{2}\nmid d$, $q_{1}\nmid d$ and $q_{2}\nmid d$ for all $i=1,2,\dots,r-2$ such that
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$
Define
$a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$
Since
$\frac{a}{(a,n)}=p_{1}p_{2}\quad\mbox{and}\quad\frac{b}{(b,n)}=q_{1}q_{2},$
therefore Lemma 2.5 implies that $a>n\ \mbox{or}\ b>n.$ Hence,
$f(\grave{u})\geq a>n\quad\mbox{or}\quad f(\grave{v})\geq b>n,$
which is absurd. The cases $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i\geq 3$ can be treated in a very similar way of case $3$ which are absurd. Hence the result follows. ∎
Example 4.3.
In Example 4.2, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$. Then, applying Theorem 4.5, we have $\omega\left(7\right)=\omega\left(8\right)=1$ and $\left\lfloor\frac{20}{7}\right\rfloor=\left\lfloor\frac{20}{8}\right\rfloor=2.$ Hence, $\deg(u)=\deg(v)=18$, (by applying Lemma 4.1).
The converse of Theorem 4.5 does not hold. For instance, let $u,v\in V(D_{23})$ such that $f(u)=10$, $f(v)=14$. By Theorem 4.2, we have $deg(u)=deg(v)=10$. However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$, which demonstrates that the equality of degrees does not imply the equality of values of the floor function.
Corollary 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f^{*}(u)>1$, $f^{*}(v)>1$. If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
Corollary 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$. If $t\mid n$ and $(t,f(u))=1$, then $\deg(u)=\deg(v)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $t\mid n$ and $(t,f(u))=1$. The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
$f^{*}(v)=\frac{f(v)}{(f(v),n)}=\frac{tf(u)}{(tf(u),n)}=\frac{tf(u)}{t(f(u),n)}=\frac{f(u)}{(f(u),n)}=f^{*}(u).$
Therefore, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same primes factors or each of them is equal to 1. Hence, using Corollary 4.3 and Theorem 4.1, the result follows. ∎
The following two examples shows the converse of Corollary 4.3 is not true,
Example 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\neq f(v)$, $f(u)\mid f(v)$, and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number, using Theorem 4.1, we have
$\deg(u)=n-1=\deg(v).$
However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
Example 4.5.
Suppose $D_{n}$ is the maximal Diophantine graph of even order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$, using Lemma 4.1 and Lemmas 2.2 part $(i)$, we have
$\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=n-2=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor=\deg(v).$
However, $f^{*}(u)=p$ and $f^{*}(v)=q$ have distinct prime factors.
The last two examples motivate the following partial converse of Corollary 4.3.
Theorem 4.6.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number and $f^{*}(u)>1$. If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$. Then, using Lemmas 2.2 part $(ii)$, we have $f^{*}(u)\mid f^{*}(v)$. Consequently, $f^{*}(v)>1$ and also any prime factor of $f^{*}(u)$ is a prime factor of $f^{*}(v)$. Let there exists a prime number $p$ such that $p\mid f^{*}(v)$. Then, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$. Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is not a prime power number, we get $s\neq v$. Using Lemma 3.1, we have $(f(v),f(s))\nmid n$. Therefore, $s\notin N(v)$. Consequently, using Lemma 4.3, we obtain $s\notin N(u)$. Then we get $\left(f(u),p^{\acute{v}_{p}(n)}\right)\nmid n$. Thus, $p^{\acute{v}_{p}(n)}\mid f(u)$. Consequently, using Lemmas 2.2 part $(i)$, $p\mid f^{*}(u)$. Therefore, for every prime number $p$,
$p\mid f^{*}(u)\quad\mbox{if and only if}\quad p\mid f^{*}(v).$
Hence, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same prime factors. ∎
Future Work
Give two efficient algorithms that one determines the maximal Diophantine graphs and anther determines whether a given graph is Diophantine or not. Furthermore, exploring different variants of Diophantine labeling can further enrich the theory and uncover new connections between graph theory and number theory.
Declaration
The authors contributed equally to the manuscripts and approved the submitted scripts. We declare that we have no conflicts of interests. We confirm that this work is original and has not been published elsewhere, nor is it currently under consideration for publication elsewhere.